4 Question6 LO1 How much would you have to invest in years 1 and 2 in order to have $5,000 in 4 years if your account earns 2% compounded annually? O $2.379.13 O $2,33248 o $2,450.50 O $2.426.71 Question 7 4 pts LO1 How much would y 5.00% APR, ou have to invest today in an account that pays compounded monthly, to have $3,000 in 4 months?
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

5000=P*(1.02)^3+P*(1.02)^2[where P=investment made in year 1 and 2]

5000=P[(1.02)^3+(1.02)^2]

5000=P*2.101608

5000/2.101608=P

Hence P=investment=$2379.13(Approx).

7.

We use the formula:
A=P(1+r/1200)^12n
where
A=future value
P=present value
r=rate of interest
n=time period.

3000=P(1+0.05/12)^4

P=3000/(1+0.05/12)^4

=(3000*0.983505508)

=$2950.52(Approx).